OpenStudy (anonymous):

The variables x and y satisfy the equation (x)^ny=C, where n and C are constants. When x=1.10, y=5.20, and when x=3.20, y=1.05. (i) Find the values of n and C.

OpenStudy (anonymous):

\[x ^{n}y=C\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

if we substitute your data, we get two different conditions, namely: \[\Large \begin{gathered} {\left( {1.1} \right)^n} \cdot 5.2 = C \hfill \\ \hfill \\ {\left( {3.2} \right)^n} \cdot 1.05 = C \hfill \\ \end{gathered} \] those equation are an algebraic system, which can be solved for n and C

OpenStudy (michele_laino):

equations*

OpenStudy (anonymous):

not being able to solve -_-

OpenStudy (michele_laino):

if I use the elimination method, I can write this: \[\Large {\left( {1.1} \right)^n} \cdot 5.2 = {\left( {3.2} \right)^n} \cdot 1.05\]

OpenStudy (anonymous):

yesss

OpenStudy (michele_laino):

now, I divide both sides of that equation by (1.1)^n, so I get: \[\Large 5.2 = {\left( {\frac{{3.2}}{{1.1}}} \right)^n} \cdot 1.05\]

OpenStudy (michele_laino):

then I divide both sides again by 1.05, so I can write this: \[\Large \frac{{5.2}}{{1.05}} = {\left( {\frac{{3.2}}{{1.1}}} \right)^n}\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

we got an exponential equation, which can be solved using logarithms

OpenStudy (anonymous):

not getting the right answer

OpenStudy (michele_laino):

why?

OpenStudy (anonymous):

i dont know.. can you continue further ?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

if we take the logarithm in base 10, of both sides, we get: \[\Large \begin{gathered} n \cdot {\log _{10}}\left( {\frac{{3.2}}{{1.1}}} \right) = {\log _{10}}\left( {\frac{{5.2}}{{1.05}}} \right) \hfill \\ \hfill \\ n = \frac{{{{\log }_{10}}\left( {\frac{{5.2}}{{1.05}}} \right)}}{{{{\log }_{10}}\left( {\frac{{3.2}}{{1.1}}} \right)}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

finally got the answer.. thank you very much

OpenStudy (michele_laino):

thanks! :)